Studies sows that CNG use for Sedan car sold in India is normally distributed with a
mean of 25.5km/ kg and a standard deviation of 4.5km/g. To get a competitive edge, a
manufacturer wishes to develop a Sedan car that outperforms 95% of the existing sedan
in fuel economy, what must be the CNG use rate for the new car be ?
1
Expert's answer
2020-08-03T18:35:39-0400

Let "X" be  the CNG use rate: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=25.5, \sigma=4.5"


"P(X<x)=0.95"

"P(Z<z_c)=0.95=>z_c=1.645"

"x=\mu+z_c\times\sigma"

"x=25.5+1.645\times4.5=32.9025\approx32.9"

The CNG use rate for the new car must be 32.9 km/ kg